# 1]FAIR at Meta

### v1

@@ 3.3 Moving past algorithm imitation via search dynamics bootstrapping
To reduce the number of search steps generated by a search-augmented model during inference, we implement a method to shift the distribution with which the decoder generates execution traces. First, a search-augmented model is trained to imitate the search dynamic of our non-deterministic $A^{*}$ implementation. This non-deterministic $A^{*}$ implementation breaks cost ties randomly and randomizes the order with which child nodes are expanded. By doing so, we do not decrease the efficiency of $A^{*}$ search itself and merely change the order with which different nodes are searched while still respecting $A^{*}$’s heuristic and cost calculations.
Using a non-deterministic implementation randomizes the sequence generation process and induces additional variance into the length of each execution trace. The resulting search-augmented model will then approximate the probability distribution with which the training sequences were generated.
Once a model is trained to imitate the search dynamics of non-deterministic $A^{*}$ search, it is used to generate a new training dataset consisting of shorter token sequences. This new dataset is constructed by using the trained search-augmented model to generate multiple token sequences for each training prompt. In this step, we only use the training dataset for bootstrapping and not the test dataset. Each generated sequence is parsed and checked if it ends in an optimal plan. If this is the case and the sequence is also shorter than the corresponding sequence contained in the original training dataset, then this shortened sequence is included in the new short sequence training dataset. If the generated sequence does not end in an optimal plan or is longer, then the sequence from the original training dataset is re-used.
Subsequently, the search-augmented model is fine-tuned on the new short sequence training dataset. To distinguish from the search-augmented model trained on the original dataset, we call this new model Searchformer. This procedure can then be repeated by using the resulting fine-tuned model to generate the next even shorter sequence dataset and then using this dataset to further fine-tune the Searchformer model. In Section 4.4 we demonstrate that this procedure does in fact reduce the number of search steps performed during inference while further improving performance. At this point, the Searchformer model no longer imitates $A^{*}$ search and has instead discovered a new way of solving a planning problem using fewer search steps.
@@ 4 Experiments
In our experiments, we use a deterministic and a non-deterministic variant of $A^{*}$ search to generate each sequence dataset.
Deterministic $A^{*}$ dataset
In the first case, sequences are generated by executing $A^{*}$ in a deterministic fashion (by ordering child nodes and breaking equal cost ties in the frontier set deterministically). Consequently, given a task prompt, the optimal plan and $A^{*}$ execution trace is unique. A Transformer model trained on this data only has to learn which token to predict next. Evaluating such a model is simple, because the generated token sequences need to exactly match the token sequences generated by $A^{*}$.
Non-deterministic $A^{*}$ dataset
In practice, there can be multiple plans that solve the same task optimally and a deterministic implementation of $A^{*}$ search will only compute one optimal solution. To diversify the dataset and include different search sequences, we generate sequences by executing $A^{*}$ search in a non-deterministic fashion: in each run, we randomly shuffle the order with which child nodes are expanded and break equal cost ties randomly. Consequently, the generated <trace><plan>-formatted sequences vary between different executions. The resulting plans may differ as well but are still optimal. This modification does not decrease the efficiency of $A^{*}$ search, but merely forces our implementation to use different search sequences that still correspond to executing the pseudo-code listed in Figure 1(c). Other randomization strategies—which do influence $A^{*}$’s efficiency—are also possible (e.g., randomizing cost values at each step) and are left for future works.
Statistics
Figure 3 presents an overview of the token sequence length used for each model type for the Sokoban dataset and each maze datasets. The length of the generated $A^{*}$ execution traces grows with the grid size of each maze. In Sokoban, the task with the largest state space, execution traces are the longest. However, the solution plans for Sokoban are amongst the shortest.
4.1 Evaluation Criteria
For each task, one model may generate a search sequence ending either in an optimal plan, a correct but sub-optimal plan, or an invalid plan. For a plan to be valid, we not only check whether the final solution sequence is feasible, but also verify the output format. The search-augmented models must generate a token sequence of the format <trace><plan>, while solution-only models must generate a token sequence of the format <plan> only. Generated sequences violating such constraints—including sequences that do not predict an end-of-sequence token eos—are omitted.
4.1.1 Measuring plan optimality
  • Any-optimal-64 criterion. For each task, we sample 64 responses from a trained model. They are parsed and then evaluated to test if the final plan is feasible and optimal, regardless of the generated <trace> part. If any of the 64 plans is feasible and optimal, then the task is labelled as correct.
    $\text{SWC}:=\frac{1}{n}\sum_{i=1}^{n}c_{i}\frac{l_{i}^{*}}{\max\{l_{i},l_{i}^{*}\}}.$
4.1.2 Measuring search dynamics length
Given the plan is correct or optimal, we evaluate the length of sequence dynamics in terms of number of tokens using one of the two criteria:
    $\text{ILR}:=\frac{1}{n}\sum_{i=1}^{n}c_{i}\frac{t_{i}^{*}}{t_{i}}.$
Unless indicated otherwise, each experiment is repeated five times and each figure plots averages across all repeats. Error bars are used to indicate the Standard Error of Measurement (SEM).
@@ 4.2 Maze Navigation
(a) Deterministic Prediction
(b) Non-deterministic Prediction
Figure 4: Predicting intermediate computation steps leads to robust performance in the low data regime. Each model was trained to predict optimal plans for $30\times 30$ maze navigation tasks. The decoder of the search-augmented models were trained on sequences of the format <trace><plan> whereas the decoder of the solution-only models were trained on sequence of the format <plan>. For each model, the number of free parameters (indicated in millions of parameters with “15M”, “46M”, and “175M”) is varied by changing the hyper-parameters of the encoder-decoder architecture. Appendix A.2 describes the used architecture and hyper-parameters in more detail.
@@
In the first experiment set, we train a set of encoder-decoder Transformer models to predict optimal plans for $30\times 30$ mazes. We vary the training dataset size and model size (the number of optimized parameters) between different training runs and evaluate each model on the test tasks generated using the same hyper-parameters (e.g., maze size).
Evaluation with the exact-match criterion
@@
Figure 4(a) plots for how many test tasks a correct response was generated. Both solution-only and search-augmented models are trained on the deterministic $A^{*}$ dataset and follow the exact-match criterion. One can observe that the solution-only model is outperformed by most search-augmented models. Only for large enough training datasets, the solution-only model matches the performance of the worst search-augmented model. In the low training data regime (100,000 training sequences and less), performance of the solution-only model degrades significantly, while the performance of each search-augmented model stays relatively high.
@@
This result is surprising, because the search-augmented models must generate a sequence of the format <trace><plan>, which is on average more than 10 times longer than the corresponding sequence of the format <plan> generated by the solution-only models. Furthermore, even the smallest search-augmented model with 15 million (15M) parameters significantly outperforms the much larger solution-only model with 175 million (175M) parameters.
Evaluation with the any-optimal-64 criterion
When trained on non-deterministic $A^{*}$ data, there could be multiple different optimal paths for one task, and in practice one would only use the model to compute an plan, the any-optimal-64 metric is more appropriate. Note that using the any-optimal-64 criterion results in higher absolute numbers compared to exact-match, because to score one point under this criterion only one generated response needs to be correct.
Figure 6: Performance across different task difficulties. Increasing task difficulty from left to right degrades performance of the solution-only models significantly. Here, each model was trained on 50,000 sequences.
@@
Figure 4(b) plots the any-optimal-64 criterion on the test set as a function of the number of non-deterministic $A^{*}$ training sequences. We can observe a similar pattern: even the smallest search-augmented models outperform solution-only model, especially for small training set. Moreover, we found that model size only impacts the performance of each of the search-augmented models when using very small training datasets (50,000 training sequences). For larger training dataset sizes no significant difference is found. Increasing the number of parameters of the solution-only models does not significantly improve their performance in the low data regime (Figure 6).
@@
Figure 6 illustrates how a task’s difficulty influences the performance of each model. Here, we focus again on the dataset generated by non-deterministic $A^{*}$, and consider the number of correctly solved test tasks as a function of maze size. The larger the maze, the larger the task’s state space and the more computation is required to find an optimal solution plan. While the solution-only model reaches close to the performance of search-augmented model for the small $10\times 10$ mazes, its performance drops rapidly as the task becomes more challenging. In contrast, the search-augmented models maintain their high accuracy as the task becomes harder, even for its smallest model size. Appendix A.5 presents A full comparison across all maze sizes.
Figure 7: Performance comparison with the Sokoban dataset. Percentage of correctly answered test prompts for each model size. All models see 100,000 training sequences. The plot averages percentage values across three different seeds and evaluates each model on the same 200 test tasks. Test tasks are not included in the training dataset.
Overall, while the solution-only models learn to predict an optimal plan if the used training dataset is large and diverse enough, search-augmented models perform significantly better in the low data regime and scale better to more difficult tasks. The search-augmented models reach higher performance because they can perform on-demand computation during inference. More specifically, the search-augmented models can control the search dynamics for a grounded reasoning chain that leads to an optimal plan, while the solution-only models have to infer direct correlations between a task description and its optimal plan through supervised learning where many of such correlations can be spurious and unreliable in the test time.
@@ 4.3 Solving Sokoban puzzles
To test if similar results can be obtained on a different and more complex task with a different tokenization pattern, we generate a planning dataset for Sokoban puzzles. Here, $A^{*}$’s execution traces are generated non-deterministically and expressed as a token sequence as outlined in Figure 10.
Figure 7 plots how often each model generated a correct optimal plan for each test task. As before, by training on execution traces, the search-augmented models outperform the solution-only models. Even increasing the parameterization of a solution-only model to 747 million parameters only leads to a marginal performance improvement. On average, this 747 million parameter solution-only model is still outperformed slightly by a smaller 175 million parameter search-augmented model. This experiment confirms the findings found for mazes on a dataset with more complex planning tasks that use a different tokenization method.
@@ 4.4 Searchformer: Improving search dynamics via bootstrapping
In this last experiment, we investigate how the search-augmented model can be iteratively improved to compute an optimal plan while predicting fewer search steps. Here, our goal is to shorten the length of the search trace while still producing an optimal solution.
We start out with the smallest search-augmented model trained on the non-deterministic $A^{*}$ Sokoban dataset and use it to generate a new shorter sequence training dataset as outlined in Sec. 3.3. For each Sokoban puzzle in the training data, we generated 32 different <trace><plan> sequences by sampling tokens from the Transformer’s output distribution and include the shortest generation (measured in tokens) if it contains an optimal plan.
Subsequently, we fine-tune the search-augmented model on this newly created training data (by running an additional 10,000 training steps) to obtain the first Searchformer model. Using this Searchformer model, we subsequently generate another short sequence dataset and repeat the fine-tuning procedure. This newly generated dataset is then used to further fine-tune the previously obtained Searchformer model for another 10,000 training steps. In total, we repeat this improvement step three times.
(a) Comparison of distribution of search dynamics length between $A^{*}$ search and our models. Each test task contributes one data point of averaged length on trials that lead to optimal plans (i.e., average-on-optimal length (Sec. 4.1.2)).
@@
(c) Distribution of lengths of individual sequences that lead to an optimal plan.
Figure 8: Improvement of search dynamics length on optimal plans via bootstrapping in Sokoban. After three improvement steps, our Searchformer model reduced the generated search dynamics token length to an average of 3094—a 26.8% reduction in total from 4222, the average $A^{*}$ search length. With each improvement step, the Searchformer models generated shorter and shorter token sequences when computing an optimal plan. 8(a): Comparison of search dynamics length (in terms of number of tokens) between $A^{*}$ search and our models (search-augmented in blue, Searchformer step 1-3 in orange), on the test subset in which our models yield optimal plans. Here, for each test task, we average the lengths of sequences that ends in an optimal plan, out of 64 trials (i.e., average-on-optimal length (Sec. 4.1.2)). The box plots show the distribution of average-on-optimal lengths, in which the left boundary, mid solid line, right boundary represents the 25%, 50% and 75% quantiles. Dotted lines are means and whiskers show the range. 8(b): Histograms of average-on-optimal lengths. 8(c): Histograms of the length of individual sequences that lead to optimal plan. Because it is more likely for each model to generate an optimal plan for easy tasks with fewer search steps, the distributions are skewed towards the left—towards shorter sequence lengths. The data plotted in this figure aggregates sequence lengths across three experiment repeats.
@@
Figure 8 illustrates how the sequence lengths generated by the Searchformer model’s are iteratively shortened by our bootstrapping method. With every improvement step, the length of the generated traces—the number of search steps—decreases (Figure 8(a)). When computing an optimal plan, the final Searchformer model generates search dynamics sequences that are on average 26.8% shorter than the sequences generated with $A^{*}$ search. Consequently, the Searchformer model found a more efficient way to plan a complicated task and go beyond $A^{*}$’s search, even for unseen test puzzles. In Figure 8(b) we can observe that the search-augmented model generates sequences that on average match the sequences generated with $A^{*}$ search in length. The Searchformer models generate shorter sequences resulting in a distribution that is skewed towards shorter sequence lengths.
We find that on average each model produces an optimal plan more often for tasks that can be solved with fewer search steps. Generating an optimal plan for more complex tasks is possible with each model but requires the model to be prompted more often. This is illustrated in Figure 8(c) which plots the sequence length distribution for individual sequences. The sequence length distribution of both models is skewed towards shorter sequences and the models do not match the $A^{*}$ sequence length distribution.
Table 1: Test set performance in the Sokoban tasks. Over 200 unseen test Sokoban tasks, we report percentage of solved, percentage of optimal solutions, Success weighted by cost (SWC) and Improved Length Ratio (ILR) w.r.t $A^{*}$ reference, averaged over sequences ending in an optimal/correct plan. Please check Sec. 4.1 for more detailed definitions of optimality metrics (Sec. 4.1.1) and search length metrics (Sec. 4.1.2). The result shows that continue fine-tuning Searchformer yields shorter search dynamics length and at the same time improved percentage of solved/optimally solved task. All values are averages with errors indicating SEM.
Params. Model Solved (%) Optimal (%) SWC ILR-on-solved ILR-on-optimal
45M Solution only 90.3 $\pm 1.0$ 86.8 $\pm 0.3$ 0.890 $\pm 0.009$
Search augmented 92.5 $\pm 1.0$ 90.8 $\pm 1.6$ 0.924 $\pm 0.011$ 0.908 $\pm 0.020$ 0.919 $\pm 0.019$
Searchformer, step 1 95.5 $\pm 1.0$ 93.5 $\pm 1.0$ 0.953 $\pm 0.010$ 1.054 $\pm 0.025$ 1.062 $\pm 0.015$
Searchformer, step 2 96.0 $\pm 0.5$ 93.4 $\pm 0.6$ 0.957 $\pm 0.005$ 1.158 $\pm 0.025$ 1.181 $\pm 0.012$
Searchformer, step 3 95.5 $\pm 0.8$ 93.7 $\pm 1.6$ 0.953 $\pm 0.009$ 1.292 $\pm 0.044$ 1.343 $\pm 0.067$
175M Solution only 95.7 $\pm 0.2$ 90.0 $\pm 0.8$ 0.949 $\pm 0.003$
Search augmented 95.2 $\pm 0.9$ 93.2 $\pm 1.0$ 0.949 $\pm 0.010$ 0.925 $\pm 0.010$ 0.933 $\pm 0.011$
757M Solution only 96.5 $\pm 0.1$ 92.2 $\pm 1.2$ 0.958 $\pm 0.002$
As reported in Table 1, fine-tuning the model resulted in a significant performance improvement, reducing the rate of incorrect and non-optimal solutions by 40% and 30% respectively. After three steps of fine-tuning, the 45M Searchformer shows similar to a 757M model that learns to predict a plan (task solution) only. The comparison using the more fine-grained SWC measures, which also take sub-optimal plans into consideration, shows a similar trend. See Sec. 4.1 for definitions of evaluation measures.
Reduced length of search dynamics with Searchformer
Lastly, we evaluate to which extend the search-augmented and Searchformer models reduced the number of generated search steps to compute a correct plan. This analysis is different from the results presented in Figure 8 because here we focus on computing length ratios between traces generated with each model and traces generated with $A^{*}$ search using Improved Length Ratio (ILR) measure.
This improvement can be observed in Table 1 for each Searchformer model where with every improvement step the model generates shorter and shorter execution traces leading to higher ILR scores. Furthermore, the scores for the two search-augmented models are slightly below one, indicating that the search-augmented models tend to perform slightly more search steps than $A^{*}$ search, while fine-tuned models has a ILR score higher than 1, showing more efficient search dynamics, in particular for those that lead to optimal plans. For example, $A^{*}$ search dynamics is $\sim 34.3\%$ longer than that produced by Searchformer after 3 steps of fine-tuning. More details about model’s performance characteristics are presented in Appendix A.4.
5 Conclusion, limitations, and future work

### v2

@@ 3.3 Moving past algorithm imitation via search dynamics bootstrapping
To reduce the number of tokens generated by a search-augmented model during inference, we implement a method to shift the distribution with which the decoder generates execution traces. First, a search-augmented model is trained to imitate the search dynamics of $A^{*}$ search. To discover new search dynamics with this search-augmented model and explore the execution trace space, the search-augmented model must generate different sequences for the same task prompt. We accomplish this by inducing non-determinism into the training data and use a non-determinsitic $A^{*}$ implementation that breaks cost ties randomly and randomizes the order with which child nodes are expanded. This approach does not decrease the efficiency of $A^{*}$ search itself and merely changes the order with which different nodes are searched while still respecting $A^{*}$’s heuristic and cost calculations. The resulting search-augmented model will then approximate the probability distribution with which the training sequences were generated.
Once a model is trained to imitate the search dynamics of non-deterministic $A^{*}$ search, it is used to generate a new training dataset consisting of shorter token sequences. This new dataset is constructed by using the trained search-augmented model to sample multiple different token sequences for each training prompt. In this step, we only use the training dataset for bootstrapping and not the test dataset. Each generated sequence is parsed and checked if it ends in an optimal plan. If this is the case and the sequence is also shorter than the corresponding sequence contained in the original training dataset, then this shortened sequence is included in the new short sequence training dataset. If the generated sequence does not end in an optimal plan or is longer than the original training sequence, then the sequence from the original training dataset is re-used.
Subsequently, the search-augmented model is fine-tuned on the new short sequence training dataset. To distinguish from the search-augmented model that imitates $A^{*}$’s search dynamics, we call this new model Searchformer. This procedure can then be repeated by using the resulting fine-tuned model to generate the next even shorter sequence dataset and then fine-tuning the Searchformer model again. In Section 4.3 we demonstrate that this procedure does in fact reduce the number of steps performed during inference while further improving performance. The Searchformer model no longer imitates $A^{*}$ search and has instead discovered a new way of solving a planning problem using fewer steps.
@@ 4 Experiments
In our experiments, we use two different $A^{*}$ implementations for sequence data generation:
  1. 1.
    Deterministic $A^{*}$ dataset: Sequences are generated by executing $A^{*}$ in a deterministic fashion (by ordering child nodes and breaking equal cost ties deterministically). Consequently, given a task prompt, the optimal plan and $A^{*}$ execution trace is unique. Here, the Transformer learns the deterministic breaking rules implicitly encoded in the data. Evaluating such a model is simple, because the generated sequences need to exactly match the sequences generated by $A^{*}$.
  2. 2.
    Non-deterministic $A^{*}$ dataset: Sequences are generated by executing $A^{*}$ in a non-deterministic fashion (by randomly ordering child nodes and breaking equal cost ties randomly). Consequently, given a task prompt, the optimal plan and $A^{*}$ execution trace is no longer unique and there are multiple correct responses. Here, the Transformer learns to generate the random tie breaking rules implicitly encoded in the sequence data. Consequently, the generated sequences vary between different executions, but the resulting plans are still optimal and execution traces still respect $A^{*}$’s cost and heuristic calculations as described in Section 3.3.
Figure 7 in Appendix C presents an overview of the token sequence length for each dataset and shows that the generated $A^{*}$ execution traces grow in length with task complexity. Figure 8 shows that training and test sets are matched in difficulty and have comparable trace lengths. For each task, one model may generate a search sequence ending either in an optimal plan, a feasible plan (a plan that is correct but sub-optimal), or an invalid plan. In Appendix D we outline how each model’s ability to predict a feasible and optimal plan is scored and details about how the search dynamics of the search-augmented and Searchformer models is evaluated.
Unless indicated otherwise, each experiment is repeated five times and each figure plots averages across all repeats. All reported errors indicate the Standard Error of Measurement (SEM).
@@ 4.1 Maze navigation
In the first experiment set, we train a set of encoder-decoder Transformer models to predict optimal plans for maze navigation tasks. We vary the training dataset size and model size (the number of optimized parameters) between different training runs and evaluate each model on the test tasks generated using the same hyper-parameters.
(a) Deterministic case
(b) Non-deterministic case
(c) Performance across task difficulties
Figure 2: Predicting intermediate computation steps leads to robust performance in the low data regime. For each model, the number of free parameters (indicated in millions of parameters with “15M”, “46M”, and “175M”) is varied. 2(a): Comparison of how many test tasks were answered with a correct token sequence when training on the deterministic $A^{*}$ dataset (exact-match criterion in Appendix D). 2(b): Comparison for how many test task at least one optimal plan was found when training on the non-deterministic $A^{*}$ dataset (any-optimal-64 criterion in Appendix D). 2(c): Performance degradation when increasing task difficulty (maze size). Here, the non-deterministic $A^{*}$ dataset was used and models are evaluated with the any-optimal-64 criterion.
Deterministic $A^{*}$
@@
Figure 2(a) plots for how many test tasks a correct response was generated. Both solution-only and search-augmented models are trained on the deterministic $A^{*}$ dataset and are evaluated if they exactly re-produce the token sequences generated with $A^{*}$ search (please refer to the exact-match criterion in Appendix D). One can observe that the solution-only model is outperformed by most search-augmented models. Only for large enough training datasets, the solution-only model matches the performance of the worst search-augmented model. In the low training data regime (100,000 training sequences and less), performance of the solution-only model degrades significantly, while the performance of each search-augmented model stays relatively high.
@@
This result is surprising, because for more than 90% of the test mazes, the search-augmented models generate <trace><plan>-formatted sequences that are thousands of tokens long without predicting any single token incorrectly. Moreover, the solution-only models, that on average predict ten times shorter sequences, are significantly outperformed by the search-augmented models. Even the smallest search-augmented model significantly outperforms the much larger solution-only model parameters.
This result highlights the power of training Transformers to generate long algorithm execution traces. We do not observe compounding prediction errors that usually limit deep model-based RL agents (Asadi et al., 2018), because the used backward-causal decoder network constructs for an $n$-step sequence an $n\times n$ attention map. Here, this property of the Transformer architecture is used to boost performance when predicting an optimal plan.
Non-deterministic $A^{*}$
When trained on non-deterministic $A^{*}$ data, the model could output multiple different optimal paths for one task. Here, we use each model to generate 64 token sequences for each task. The test task is counted as correctly answered of any one of the 64 sequences contains an optimal plan (please refer to the any-optimal-64 criterion in Appendix D). Because we only test if at least one generated sequence contains an optimal plan, we obtain higher absolute numbers in Figure 2(b) than in Figure 2(a).
@@
Figure 2(b) plots for how many test tasks an optimal plan was found when generating for each test task 64 token sequences. Here, we can observe a pattern similar to the deterministic $A^{*}$ dataset: even the smallest search-augmented models outperform solution-only model, especially for a small training set. Moreover, we found that model size only impacts the performance of each of the search-augmented models when using very small training datasets (50,000 training sequences). For larger training dataset sizes no significant difference is found. Increasing the number of parameters of the solution-only models does not significantly improve their performance in the low data regime (Figure 9 in Appendix F).
@@
Lastly, Figure 2(c) illustrates how a task’s difficulty influences the performance of each model. Here, we focus again on the dataset generated by non-deterministic $A^{*}$, and consider the number of correctly solved test tasks as a function of maze size. The larger the maze, the larger the task’s state space and the more computation is required to find an optimal solution plan. While the solution-only model’s performance drops rapidly as the tasks become more challenging, the search-augmented models maintain a comparably high accuracy, even for its smallest model size. Appendix F presents a full comparison across all maze sizes.
Overall, while the solution-only models learn to predict an optimal plan if the used training dataset is large and diverse enough, search-augmented models perform significantly better in the low data regime and scale better to more difficult tasks. The search-augmented models reach higher performance because they can perform on-demand computation during inference. More specifically, the search-augmented models imitate the search dynamics for a grounded reasoning chain that leads to an optimal plan, while the solution-only models have to infer direct correlations between a task description and its optimal plan through supervised learning where many of such correlations can be spurious and unreliable during evaluation on the test task set.
@@ 4.2 Solving Sokoban puzzles
Table 1: Test set performance in the Sokoban tasks. Over 200 unseen test Sokoban tasks, we report percentage of solved and optimally solved test tasks. For sequences ending in either an optimal and correct plan we report the SWC (Success Weighted by Cost) score, and ILR (Improved Length Ratio of Search Dynamics) scores. The better trace and solution quality, the higher the scores. Please check Appendix D for detailed definitions of these scores.
Params. Model Solved (%) Optimal (%) SWC ($\uparrow$) ILR (solved, $\uparrow$) ILR (optimal, $\uparrow$)
45M Solution only 90.3 ±$\pm 1.0$ 86.8 ±$\pm 0.3$ 0.890 ±$\pm 0.009$
Search augmented 92.5 ±$\pm 1.0$ 90.8 ±$\pm 1.6$ 0.924 ±$\pm 0.011$ 0.908 ±$\pm 0.020$ 0.919 ±$\pm 0.019$
Searchformer, step 1 95.5 ±$\pm 1.0$ 93.5 ±$\pm 1.0$ 0.953 ±$\pm 0.010$ 1.054 ±$\pm 0.025$ 1.062 ±$\pm 0.015$
Searchformer, step 2 96.0 ±$\pm 0.5$ 93.4 ±$\pm 0.6$ 0.957 ±$\pm 0.005$ 1.158 ±$\pm 0.025$ 1.181 ±$\pm 0.012$
Searchformer, step 3 95.5 ±$\pm 0.8$ 93.7 ±$\pm 1.6$ 0.953 ±$\pm 0.009$ 1.292 ±$\pm 0.044$ 1.343 ±$\pm 0.067$
175M Solution only 95.7 ±$\pm 0.2$ 90.0 ±$\pm 0.8$ 0.949 ±$\pm 0.003$
Search augmented 95.2 ±$\pm 0.9$ 93.2 ±$\pm 1.0$ 0.949 ±$\pm 0.010$ 0.925 ±$\pm 0.010$ 0.933 ±$\pm 0.011$
757M Solution only 96.5 ±$\pm 0.1$ 92.2 ±$\pm 1.2$ 0.958 ±$\pm 0.002$
To test if similar results can be obtained on a different and more complex task with a different tokenization pattern and different transition structure, we repeat our experiments for Sokoban puzzles using our non-deterministic $A^{*}$ implementation. Table 1 lists how often each model generated a correct optimal plan for each test task. As before, by training on execution traces, the search-augmented models outperform the solution-only models. Even increasing the parameterization of a solution-only model to 747 million parameters only leads to a marginal performance improvement. On average, this 747 million parameter solution-only model is still outperformed slightly by a smaller 175 million parameter search-augmented model. This experiment further confirms our findings on more complex planning tasks with a different transition structure and a different tokenization method.
@@ 4.3 Searchformer: Improving search dynamics via bootstrapping
In this last experiment, we investigate how the search-augmented model can be iteratively improved to compute an optimal plan while generating a shorter execution trace. Here, our goal is to shorten the length of the search trace while still producing an optimal solution.
We start out with the smallest search-augmented model trained on the non-deterministic $A^{*}$ Sokoban dataset and use it to generate a new shorter sequence training dataset as outlined in Section 3.3. For each Sokoban puzzle in the training data, we generated 32 different <trace><plan>-formatted sequences by sampling tokens from the Transformer’s output distribution and include the shortest generation (measured in tokens) if it contains an optimal plan. Subsequently, we fine-tune the search-augmented model on this newly created training data (by running an additional 10,000 training steps) to obtain the first Searchformer model. Using this Searchformer model, we subsequently generate another short sequence dataset and repeat the fine-tuning procedure to further improve the model.
(a) Search length improvement.
@@
Figure 3: Improvement of search dynamics length via bootstrapping in Sokoban 3(a): For each Sokoban test task that was answered with an optimal plan, the average generated execution trace length is computed and averaged. The $A^{*}$ reference values are computed by generating an equal number of execution traces for each test task. For each iteration, we compare the subset of test tasks that were answered with an optimal plan by the search-augmented and Searchformer model and compare the relative improvement over our $A^{*}$ reference implementation. Figure 13 (Appendix F) shows a box plot with more details about the sequence length distribution. 3(b): Distribution of execution trace lengths for each optimally answered test task. Here, the sequence lengths were first averaged per test task and the resulting distribution is plotted in each histogram. The search-augmented model is trained to imitate our $A^{*}$ implementation. After three iterations of search dynamics bootstrapping, the resulting Searchformer model generates on average shorter execution traces.
@@
Figure 3(a) illustrates how the sequence lengths generated by the Searchformer model’s are iteratively shortened with our search dynamics boostrapping method. With every improvement step, the average length of the generated traces—the number of search steps—decreases (Figure 3(a)). When computing an optimal plan, the final Searchformer model generates search dynamics sequences that are on average 26.8% shorter than the sequences generated with $A^{*}$ search. Consequently, the Searchformer model found a way to find a plan in a complex task that is more efficient in terms of search steps than the $A^{*}$ implementation used to train the initial search-augmented model. In Figure 3(b) we can observe that the search-augmented model generates sequences that on average match the sequences generated with $A^{*}$ search in length. The Searchformer models generate shorter sequences resulting in a distribution that is skewed towards shorter sequence lengths.
As reported in Table 1, fine-tuning the model resulted in a significant performance improvement, reducing the rate of incorrect and non-optimal solutions by 40% and 30% respectively. The Success Weighted by Cost (SWC) score (Wu et al., 2019) factors in how many test tasks were solved correctly and how close the predicted plans are to the optimal length (Appendix D). Here, a perfect score would be one, and one can see in Table 1 that the comparably small Searchformer matches the performance of the largest solution-only model (also note the small SEM values). Furthermore, the Improved Length Ratio of Search Dynamics (ILR) measures how much the length of each execution trace is shortened (Appendix D). With each improvement iteration the scores increase and climb above one. For example, $A^{*}$ search dynamics is $\sim 34.3\%$ longer than the sequences generated by Searchformer after 3 steps of fine-tuning.
The results reported in Figure 3 and in Table 1 only compare each model’s performance on test tasks that were either correctly or optimally solved. To test if a model overfits only on easier test tasks with shorter execution traces, we plot in Figure 10 (in Appendix E) the execution trace length generated with $A^{*}$ against the execution trace length generated by each model as a scatter plot. Each point in this plot corresponds to a single test task. Here, the trend of shortening the execution trace via search dynamics bootstrapping is clear and one can also observe that neither model only specializes on solving easier test tasks with shorter execution traces.
5 Discussion